two cards are drwa from a well-shuffled deck of 52 cards.
what is the probability of getting two jacks?
what is the probability of getting two clubs?

Answer by nerdybill(5413)

what is the probability of getting two jacks?
since there are a total of 4 jacks in a deck of 52, the probability of the getting the first jack is 4:52 = 4/52 = 1/13
.
Now, after the first jack is removed, there are only 3 jacks left in a deck of 51 so the probability of the second card drawn is:
3:51 = 3/51 = 1/17
.
Therefore the probability is:
1/13 * 1/17 = 0.0045
.
what is the probability of getting two clubs?
Since there are a total of 13 clubs in a deck of 52, the probability of the getting the first club is 13:52 = 1/4
.
Now, after the first club is removed, there are only 12 left in a deck of 51 so the probability of the second card drawn is:
12:51 = 12/51
.
Therefore the probability is:
1/4 * 12/51
= 1/1 * 3/51
= 3/51
= 0.0588